OpenStudy (anonymous):

Find the Vertices, Foci and Asymptotes of this conic: 16y^2 -x^2 = 1

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ \cfrac{(y-k)^2}{a^2}-\cfrac{(x-h)^2}{b^2}=1}} \\ \quad \\ \quad \\ 16y^2 -x^2 = 1\implies 4^2y^2 -x^2 = 1\implies \cfrac{1}{4^{-2}}\cdot y^2 -x^2 = 1 \\ \quad \\ \cfrac{y^2}{\frac{1}{16}}-\cfrac{x^2}{1}=1\implies \cfrac{(y-0)^2}{\frac{1}{16}}-\cfrac{(x-0)^2}{1}=1 \\ \quad \\ \cfrac{(y-0)^2}{\left(\sqrt{\frac{1}{16}}\right)^2}-\cfrac{(x-0)^2}{1^2}=1\\ -----------------------------\\ vertices\implies (h\quad ,\quad k\pm a )\qquad foci\implies (h\quad ,\quad k\pm \sqrt{a^2+b^2}) \\ \quad \\ asymptotes\implies k\pm \cfrac{a}{b}(x-h)\)